what is the area of the rectangle whose length is 5 more than twice its width and its perimeter is 70 meters?
Answer by checkley77(12844)

L=2W+5
2L+2W=70
2(2W+5)+2W=70
4W+10+2W=70
6W=70-10
6W=60
W=60/6
W=10 ANS. FOR THE WIDTH.
L=2*10+5
L=20+5=25 ANS FOR THE LENGTH.
PROOF:
2*25+2*10=70
50+20=70
70=70